A light-rail train going from one station to the next on a straight section of track accelerates from rest at 2.15 m/s2 for 15.4 s/ It then proceeds at constant speed for 1180 m before slowing down at 2.2 m/s2 until it stops at the station. What is the distance between the stations?
Answer #1

Total distance =Distance covered while accelerating +1180 m +Distance covered while decelerating

Acceleration

initial velocity, u =0m/s

acceleration, a =2.15 m/s2

time, t =15.4s

Use Formula sut +1/2at2

Sj = 0m/s * 15.4s + 0.5 * 2.15m/s2 * (15.492

Sj 0 0.5 * 2.15 * 15.42

s1 254.947m

--------------------

Now find the constant speed attained

Use Formula v=u+at

v=0m/s+2.15m/s^{2}*15.4s

v-33.1lm/s

--------------------

Now distance covered while decelerating

Final velocity is zero

Use Formula 205

(Om/s)2- (33.11m/s)22.2m/s s

02-33.11°=-2 * 2.2 * s

33.11^{2}/(2*2.2)=s_{3}

s_{3}=249.153

===============

Total distance =si 1 180m + s3 = 254.947m + 1 180m + 249. 153m

ANSWER: 1684.1m

=================